# https://kaist-viclab.github.io/SLiM_site/

### 11 Mar 2026 v1

@@ 1 Introduction
Skeleton data serves as a compact yet informative modality for human action understanding, offering inherent robustness against dynamic backgrounds, lighting variations, and viewpoint shifts compared to RGB-based counterparts [33, 40, 15]. While supervised learning on skeleton data [4, 6, 9] has achieved remarkable success, it faces fundamental hurdles: beyond the prohibitive cost of large-scale annotation, supervised models often struggle with generalization, frequently overfitting to specific domain distributions and failing to adapt to unseen scenarios.
Consequently, Self-Supervised Learning (SSL) [2, 24, 42, 3, 12] has emerged as the dominant methodology to learn generalizable motion representations from abundant unlabeled data. The evolution of SSL in skeleton analysis has primarily shifted from Contrastive Learning (CL) [3] to Masked Auto-Encoder (MAE) [12] architectures. Early CL approaches [17, 39, 23, 11] focused on instance discrimination, pulling augmented views of the same sequence together. However, CL relies heavily on global pooling, which often leads to the vanishing local details problem—overlooking fine-grained motion patterns crucial for complex actions—and suffers from the difficulty of constructing semantically consistent pairs.
@@
SkeletonMAE [38] ICCV’23 MAE $30\times 25$ 28.32 1.97 17.70 74.8 77.7
MAMP [22] ICCV’23 MAE $30\times 25$ 28.32 1.97 17.70 84.9 89.1
S-JEPA [1] ECCV’24 MAE $30\times 25$ 28.32 1.97 17.70 28.32 85.3 89.8
GFP [32] ICCV’25 MAE $30\times 25$ 28.32 1.97 1.57 0.64 85.9 92.0
@@ 2.1 Contrastive learning (CL)-based Approaches
Early self-supervised approaches primarily utilized CL [3] to capture discriminative global features via instance discrimination. CrosSCLR [45] introduced a cross-view consistency constraint to learn view-invariant representations, while AimCLR [11] enhanced robustness through extreme augmentations. To incorporate richer semantics, HaLP [25] proposed latent positive mining, and ActCLR [17] decomposed actions into static and dynamic components. However, these frameworks often rely on generic spatial transformations that overlook anatomical constraints.
@@ 2.2 Masked auto-encoder (MAE)-based Approaches
MAE [12] approaches have recently shifted the focus toward reconstructing masked inputs to learn detailed skeletal-temporal dependencies. SkeletonMAE [38] established the first skeleton-based MAE framework by reconstructing masked joint coordinates to capture spatial topology. MAMP [22] advanced this by targeting motion-aware trajectories, encouraging the model to learn complex temporal dynamics. S-JEPA [1] moved beyond low-level reconstruction by employing a joint embedding predictive architecture that operates in the latent feature space. GFP [32] introduced a generative feature prediction task to bridge the gap between local reconstruction and high-level semantic understanding. While these methods have evolved from random masking [38] to motion-aware masking [22, 1, 32] strategies, they fundamentally rely on independent joint masking. Due to strong dependencies between skeletal joints, such an approach often allows models to simply interpolate missing joints from their immediate neighbors.
@@ 2.3 Other Pretext Tasks
Beyond CL and MAE, other emerging paradigms have explored generative and fusion-based strategies. MacDiff [37] and IGM [16] leverage diffusion-based generative processes to learn robust action representations by denoising skeletal sequences through iterative refinement. In terms of multi-modal integration, HSP [35] proposes a hierarchical semantic preservation framework that fuses 2D and 3D skeleton data to exploit complementary geometric and appearance cues. While these methods offer unique perspectives, they often require complex sampling or multi-stream inputs (e.g., bone or motion modalities). In contrast, SLiM achieves both superior performance and higher efficiency using a minimalist single-stream architecture focused solely on 3D joint-modality inputs.
@@ 3.1 Overview of SLiM
Fig. 2 illustrates our SLiM (Skeleton Less is More), a unified framework that seamlessly integrates masked modeling with contrastive learning without relying on a computationally heavy reconstruction decoder. The SLiM architecture adopts a teacher–student distillation scheme [34]. Specifically, the student network $f_{\bm{\theta}}$, parameterized by $\bm{\theta}$, is trained via back-propagation. The teacher network $f_{\bm{\phi}}$, parameterized by $\bm{\phi}$, is not updated by gradients but evolves via an Exponential Moving Average (EMA) [34] of the student weights: $\bm{\phi}\leftarrow\tau\bm{\phi}+(1-\tau)\bm{\theta}$, where $\tau$ is a momentum coefficient. Both networks share the identical architecture, consisting of an encoder $g$ and lightweight (3-layer MLP) prediction heads $h_{\texttt{CL}}$ and $h_{\texttt{MAE}}$.
Architecture of encoder. We employ a Vision Transformer (ViT) [8] as the backbone encoder $g$. Let the input skeleton sequence be denoted as $\mathbf{X}\in\mathbb{R}^{T\times J\times C_{\text{in}}}$, where $T$ is the sequence length, $J$ is the number of joints, and $C_{\text{in}}(=3)$ is the joint dimension. We reshape it into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{N\times D_{p}}$, where $N=(T/P_{T})\times(J/P_{J})$ is the number of patches and $D_{p}=P_{T}\cdot P_{J}\cdot C_{\text{in}}$ is the raw patch dimension. Here, $(P_{T},P_{J})$ denotes the resolution of each skeletal-temporal patch. To effectively capture the skeletal-temporal dynamics, we utilize a composite embedding strategy. First, we apply a linear projection $\mathbf{E}\in\mathbb{R}^{D_{p}\times D}$ to $\mathbf{X}_{p}$ and add absolute joint positional embeddings $\mathbf{E}_{\text{skel}}\in\mathbb{R}^{(J/P_{J})\times D}$ to encode the identity of each joint:
@@
where $\mathbf{Z}_{p}\in\mathbb{R}^{N\times D}$ represents the sequence of patch tokens and $D$ is the embedding dimension. $\mathbf{E}_{\text{skel}}^{b}\in\mathbb{R}^{N\times D}$ is a broadcast version of $\mathbf{E}_{\text{skel}}$ along the temporal dimension to match the sequence length $N$. We append a learnable class token $\mathbf{z}_{\texttt{cls}}\in\mathbb{R}^{1\times D}$ to the patch tokens $\mathbf{Z}_{p}$, initializing the input sequence $\mathbf{Z}_{0}=\left[\mathbf{z}_{\texttt{cls}}\mid\mathbf{Z}_{p}\right]$. Second, to explicitly model temporal dependencies, we adopt 1D temporal rotary positional embeddings (RoPE) [27]. RoPE is injected directly into the attention mechanism at each layer:
@@ 3.2 Decoder-Free Masked Modeling with Contrastive Learning
where $\mathbf{e}_{\texttt{[mask]}}\in\mathbb{R}^{D}$ is a learnable mask token vector, and $\mathbf{e}_{\texttt{[mask]}}^{b}\in\mathbb{R}^{N\times D}$ is broadcast version. The student then predicts the features of the masked patches: $\left[\mathbf{y}_{\text{G1},\texttt{cls}}^{s}\mid\mathbf{Y}_{\text{G1},p}^{s}\right]=f_{\bm{\theta}}\left(\left[\mathbf{z}_{\texttt{cls}}\mid\mathbf{Z}_{\text{G1},p}^{\texttt{mask}}\right]\right)$. Following the iBOT framework [42], we employ the cross-entropy loss between $f_{\bm{\theta}}$’s prediction and $f_{\bm{\phi}}$’s target distribution on the masked patches. The masked feature modeling loss is defined as:
@@
It is reminded that local views ($\mathbf{X}_{\text{L1}},\mathbf{X}_{\text{L2}},\mathbf{X}_{\text{L3}}$) are sampled from the same interval of $\mathbf{X}_{\text{G1}}$. The contrastive loss [2, 24, 5] is minimized by matching the student’s predictions for all views $\mathbf{X}_{k}\in\mathcal{V}$ to the teacher’s anchor distribution:
@@ 3.3 Semantic Tube Masking (STM)
Fig. 3-(a) and -(e) show an independent joint masking of previous methods [38, 22, 1, 32] and our masking strategy, respectively. Independent joint masking is often susceptible to trivial solutions due to the high correlation in skeleton sequences. To resolve this, we propose Semantic Tube Masking (STM), a strategy that masks anatomical joint groups (e.g., left arm, right leg, or torso) continuously along the temporal axis, thereby forming a skeletal-temporal tube. Unlike prior masking schemes that treat joints as isolated tokens, STM imposes a more challenging pretext task by occluding entire functional units of the body over time. To our knowledge, SLiM is the first to adopt tube masking for skeleton data, providing a simple yet highly effective strategy for robust representation learning.
@@ 3.4 Skeleton-Aware Augmentations (SAA)
Standard geometric augmentations [17, 11, 45] often disregard the articulated nature of the human body, leading to implausible poses or distorted limb proportions. We introduce Skeletal-Aware Augmentations (SAA), a set of improved transformations designed to construct diverse yet structurally valid pairs.
Skeleton-aware rotation. This is illustrated in Fig. 3-(f) and compared with the uniform rotation in Fig. 3-(b). Standard rotation [17, 4, 11] is typically confined to a limited range of angles. To maintain a realistic upright posture, we apply a full $360^{\circ}$ random rotation exclusively around the vertical $Y$-axis. For the remaining non-gravity axes ($X,Z$), we constrain perturbations to small angles (e.g., $30^{\circ}$) to avoid physically implausible tilting. Unlike previous works that restrict all axes to maintain stability, our approach decouples the rotation logic to maximize view diversity while preserving physical validity.
Skeleton-aware mirroring. As shown in Fig. 3-(g), and compared with the index swapping in Fig. 3-(c), heuristic index swapping [17, 11, 6] is a common approach but fails to achieve a mathematically rigorous reflection, thereby resulting in unnatural poses. Our strategy performs a geometric mirroring that simultaneously negates coordinate values along the lateral axis ($X$-axis) and reassigns joint indices according to bilateral symmetry. This dual-operation ensures that the resulting motion is anatomically consistent–preserving the correct “Front” orientation and left-right semantics–rather than producing a distorted artifact.
Bone-aware scaling. This is depicted in Fig. 3-(h) and compared with the naive scaling in Fig. 3-(d). The naive scaling of joint coordinates [11, 4] can disrupt the structure of the skeleton. To preserve the essence of the action, we decompose each bone into a directional unit vector $\hat{\mathbf{v}}$ and a scalar length $l$. We apply random scaling factors only to $l$ while strictly maintaining $\hat{\mathbf{v}}$. This effectively simulates the same action performed by subjects of varying body sizes and proportions, thereby enriching the intra-class variance without semantic distortion.
@@ 4.1 Datasets
We evaluate our SLiM on three large-scale benchmark datasets: NTU RGB+D 60 (NTU-60) [26], NTU RGB+D 120 (NTU-120) [19], and PKU-MMD II [18]. NTU-60 contains 56,880 skeleton sequences across 60 action classes, captured from three distinct camera views. It offers two evaluation protocols: Cross-Subject (X-Sub) and Cross-View (X-View). NTU-120 extends this to 120 classes and 114,480 samples, featuring Cross-Subject (X-Sub) and Cross-Setup (X-Set) protocols. PKU-MMD II encompasses 51 action classes; following standard practice, we utilize the skeleton modality under the Cross-Subject (X-Sub) protocol.
Focus on 3D skeleton. 2D skeleton sequences are typically extracted from images via pre-trained pose estimation networks [30] and inherently represent coordinates on the image plane. Consequently, they can easily achieve high performance by being back-projected onto the image space to leverage existing image-domain representations [9]. In contrast, 3D skeleton data encapsulates the pure geometric structure of human motion. Therefore, we focus our evaluation on 3D sequences to rigorously validate the model’s intrinsic geometric representation capability without any reliance on such image-domain priors.
@@ 4.2 Experiment Details
Architecture. To ensure a fair comparison with previous MAE-based methods [38, 22, 1, 32], we employ a standard ViT [8] as our backbone encoder $g$. The encoder $g$ consists of 8 layers with a hidden dimension of $D=256$ and 8 attention heads, maintaining an identical architectural complexity to existing baselines. Input skeleton sequences are fixed to a temporal length of $T=64$ frames with $J=25$ joints. For patch embedding, we adopt a skeletal-temporal patch size of $P_{J}=1$ and $P_{T}=8$.
Pre-training setup. Our SLiM is pre-trained for 150 epochs within a teacher–student distillation framework [34]. The teacher’s momentum $\tau$ is updated from 0.994 to 1. Optimization is performed using AdamW [21] with a base learning rate of $2.0\times 10^{-4}$, incorporating a 20-epoch linear warmup followed by a cosine decay [20] down to $1.0\times 10^{-6}$. We train the model with a total batch size of 768 across 4 NVIDIA RTX A6000 GPUs.

### 12 Mar 2026 v2

@@ 1 Introduction
Skeleton data serves as a compact yet informative modality for human action understanding, offering inherent robustness against dynamic backgrounds, lighting variations, and viewpoint shifts compared to RGB-based counterparts [Review1, Review2, Review3]. While supervised learning on skeleton data [CTRGCN, Skateformer, PoseC3D] has achieved remarkable success, it faces fundamental hurdles: beyond the prohibitive cost of large-scale annotation, supervised models often struggle with generalization, frequently overfitting to specific domain distributions and failing to adapt to unseen scenarios.
Consequently, Self-Supervised Learning (SSL) [DINO, DINOv2, iBOT, CL, MAE] has emerged as the dominant methodology to learn generalizable motion representations from abundant unlabeled data. The evolution of SSL in skeleton analysis has primarily shifted from Contrastive Learning (CL) [CL] to Masked Auto-Encoder (MAE) [MAE] architectures. Early CL approaches [ActCLR, CPM, CMD, AimCLR] focused on instance discrimination, pulling augmented views of the same sequence together. However, CL relies heavily on global pooling, which often leads to the vanishing local details problem—overlooking fine-grained motion patterns crucial for complex actions—and suffers from the difficulty of constructing semantically consistent pairs.
@@
SkeletonMAE [SkeletonMAE] ICCV’23 MAE $30\times 25$ 28.32 1.97 17.70 74.8 77.7
MAMP [MAMP] ICCV’23 MAE $30\times 25$ 28.32 1.97 17.70 84.9 89.1
S-JEPA [SJEPA] ECCV’24 MAE $30\times 25$ 28.32 1.97 17.70 28.32 85.3 89.8
GFP [GFP] ICCV’25 MAE $30\times 25$ 28.32 1.97 1.57 0.64 85.9 92.0
@@ 2.1 Contrastive learning (CL)-based Approaches
Early self-supervised approaches primarily utilized CL [CL] to capture discriminative global features via instance discrimination. CrosSCLR [CrosSCLR] introduced a cross-view consistency constraint to learn view-invariant representations, while AimCLR [AimCLR] enhanced robustness through extreme augmentations. To incorporate richer semantics, HaLP [HaLP] proposed latent positive mining, and ActCLR [ActCLR] decomposed actions into static and dynamic components. However, these frameworks often rely on generic spatial transformations that overlook anatomical constraints.
@@ 2.2 Masked auto-encoder (MAE)-based Approaches
MAE [MAE] approaches have recently shifted the focus toward reconstructing masked inputs to learn detailed skeletal-temporal dependencies. SkeletonMAE [SkeletonMAE] established the first skeleton-based MAE framework by reconstructing masked joint coordinates to capture spatial topology. MAMP [MAMP] advanced this by targeting motion-aware trajectories, encouraging the model to learn complex temporal dynamics. S-JEPA [SJEPA] moved beyond low-level reconstruction by employing a joint embedding predictive architecture that operates in the latent feature space. GFP [GFP] introduced a generative feature prediction task to bridge the gap between local reconstruction and high-level semantic understanding. While these methods have evolved from random masking [SkeletonMAE] to motion-aware masking [MAMP, SJEPA, GFP] strategies, they fundamentally rely on independent joint masking. Due to strong dependencies between skeletal joints, such an approach often allows models to simply interpolate missing joints from their immediate neighbors.
@@ 2.3 Other Pretext Tasks
Beyond CL and MAE, other emerging paradigms have explored generative and fusion-based strategies. MacDiff [MacDiff] and IGM [IGM] leverage diffusion-based generative processes to learn robust action representations by denoising skeletal sequences through iterative refinement. In terms of multi-modal integration, HSP [HSP] proposes a hierarchical semantic preservation framework that fuses 2D and 3D skeleton data to exploit complementary geometric and appearance cues. While these methods offer unique perspectives, they often require complex sampling or multi-stream inputs (e.g., bone or motion modalities). In contrast, SLiM achieves both superior performance and higher efficiency using a minimalist single-stream architecture focused solely on 3D joint-modality inputs.
@@ 3.1 Overview of SLiM
Fig. 2 illustrates our SLiM (Skeleton Less is More), a unified framework that seamlessly integrates masked modeling with contrastive learning without relying on a computationally heavy reconstruction decoder. The SLiM architecture adopts a teacher–student distillation scheme [Teacher]. Specifically, the student network $f_{\bm{\theta}}$, parameterized by $\bm{\theta}$, is trained via back-propagation. The teacher network $f_{\bm{\phi}}$, parameterized by $\bm{\phi}$, is not updated by gradients but evolves via an Exponential Moving Average (EMA) [Teacher] of the student weights: $\bm{\phi}\leftarrow\tau\bm{\phi}+(1-\tau)\bm{\theta}$, where $\tau$ is a momentum coefficient. Both networks share the identical architecture, consisting of an encoder $g$ and lightweight (3-layer MLP) prediction heads $h_{\texttt{CL}}$ and $h_{\texttt{MAE}}$.
Architecture of encoder. We employ a Vision Transformer (ViT) [ViT] as the backbone encoder $g$. Let the input skeleton sequence be denoted as $\mathbf{X}\in\mathbb{R}^{T\times J\times C_{\text{in}}}$, where $T$ is the sequence length, $J$ is the number of joints, and $C_{\text{in}}(=3)$ is the joint dimension. We reshape it into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{N\times D_{p}}$, where $N=(T/P_{T})\times(J/P_{J})$ is the number of patches and $D_{p}=P_{T}\cdot P_{J}\cdot C_{\text{in}}$ is the raw patch dimension. Here, $(P_{T},P_{J})$ denotes the resolution of each skeletal-temporal patch. To effectively capture the skeletal-temporal dynamics, we utilize a composite embedding strategy. First, we apply a linear projection $\mathbf{E}\in\mathbb{R}^{D_{p}\times D}$ to $\mathbf{X}_{p}$ and add absolute joint positional embeddings $\mathbf{E}_{\text{skel}}\in\mathbb{R}^{(J/P_{J})\times D}$ to encode the identity of each joint:
@@
where $\mathbf{Z}_{p}\in\mathbb{R}^{N\times D}$ represents the sequence of patch tokens and $D$ is the embedding dimension. $\mathbf{E}_{\text{skel}}^{b}\in\mathbb{R}^{N\times D}$ is a broadcast version of $\mathbf{E}_{\text{skel}}$ along the temporal dimension to match the sequence length $N$. We append a learnable class token $\mathbf{z}_{\texttt{cls}}\in\mathbb{R}^{1\times D}$ to the patch tokens $\mathbf{Z}_{p}$, initializing the input sequence $\mathbf{Z}_{0}=\left[\mathbf{z}_{\texttt{cls}}\mid\mathbf{Z}_{p}\right]$. Second, to explicitly model temporal dependencies, we adopt 1D temporal rotary positional embeddings (RoPE) [RoPE]. RoPE is injected directly into the attention mechanism at each layer:
@@ 3.2 Decoder-Free Masked Modeling with Contrastive Learning
where $\mathbf{e}_{\texttt{[mask]}}\in\mathbb{R}^{D}$ is a learnable mask token vector, and $\mathbf{e}_{\texttt{[mask]}}^{b}\in\mathbb{R}^{N\times D}$ is broadcast version. The student then predicts the features of the masked patches: $\left[\mathbf{y}_{\text{G1},\texttt{cls}}^{s}\mid\mathbf{Y}_{\text{G1},p}^{s}\right]=f_{\bm{\theta}}\left(\left[\mathbf{z}_{\texttt{cls}}\mid\mathbf{Z}_{\text{G1},p}^{\texttt{mask}}\right]\right)$. Following the iBOT framework [iBOT], we employ the cross-entropy loss between $f_{\bm{\theta}}$’s prediction and $f_{\bm{\phi}}$’s target distribution on the masked patches. The masked feature modeling loss is defined as:
@@
It is reminded that local views ($\mathbf{X}_{\text{L1}},\mathbf{X}_{\text{L2}},\mathbf{X}_{\text{L3}}$) are sampled from the same interval of $\mathbf{X}_{\text{G1}}$. The contrastive loss [DINO, DINOv2, Register] is minimized by matching the student’s predictions for all views $\mathbf{X}_{k}\in\mathcal{V}$ to the teacher’s anchor distribution:
@@ 3.3 Semantic Tube Masking (STM)
Fig. 3-(a) and -(e) show an independent joint masking of previous methods [SkeletonMAE, MAMP, SJEPA, GFP] and our masking strategy, respectively. Independent joint masking is often susceptible to trivial solutions due to the high correlation in skeleton sequences. To resolve this, we propose Semantic Tube Masking (STM), a strategy that masks anatomical joint groups (e.g., left arm, right leg, or torso) continuously along the temporal axis, thereby forming a skeletal-temporal tube. Unlike prior masking schemes that treat joints as isolated tokens, STM imposes a more challenging pretext task by occluding entire functional units of the body over time. To our knowledge, SLiM is the first to adopt tube masking for skeleton data, providing a simple yet highly effective strategy for robust representation learning.
@@ 3.4 Skeleton-Aware Augmentations (SAA)
Standard geometric augmentations [ActCLR, AimCLR, CrosSCLR] often disregard the articulated nature of the human body, leading to implausible poses or distorted limb proportions. We introduce Skeletal-Aware Augmentations (SAA), a set of improved transformations designed to construct diverse yet structurally valid pairs.
Skeleton-aware rotation. This is illustrated in Fig. 3-(f) and compared with the uniform rotation in Fig. 3-(b). Standard rotation [ActCLR, CTRGCN, AimCLR] is typically confined to a limited range of angles. To maintain a realistic upright posture, we apply a full $360^{\circ}$ random rotation exclusively around the vertical $Y$-axis. For the remaining non-gravity axes ($X,Z$), we constrain perturbations to small angles (e.g., $30^{\circ}$) to avoid physically implausible tilting. Unlike previous works that restrict all axes to maintain stability, our approach decouples the rotation logic to maximize view diversity while preserving physical validity.
Skeleton-aware mirroring. As shown in Fig. 3-(g), and compared with the index swapping in Fig. 3-(c), heuristic index swapping [ActCLR, AimCLR, Skateformer] is a common approach but fails to achieve a mathematically rigorous reflection, thereby resulting in unnatural poses. Our strategy performs a geometric mirroring that simultaneously negates coordinate values along the lateral axis ($X$-axis) and reassigns joint indices according to bilateral symmetry. This dual-operation ensures that the resulting motion is anatomically consistent–preserving the correct “Front” orientation and left-right semantics–rather than producing a distorted artifact.
Bone-aware scaling. This is depicted in Fig. 3-(h) and compared with the naive scaling in Fig. 3-(d). The naive scaling of joint coordinates [AimCLR, CTRGCN] can disrupt the structure of the skeleton. To preserve the essence of the action, we decompose each bone into a directional unit vector $\hat{\mathbf{v}}$ and a scalar length $l$. We apply random scaling factors only to $l$ while strictly maintaining $\hat{\mathbf{v}}$. This effectively simulates the same action performed by subjects of varying body sizes and proportions, thereby enriching the intra-class variance without semantic distortion.
@@ 4.1 Datasets
We evaluate our SLiM on three large-scale benchmark datasets: NTU RGB+D 60 (NTU-60) [NTU60], NTU RGB+D 120 (NTU-120) [NTU120], and PKU-MMD II [PKU]. NTU-60 contains 56,880 skeleton sequences across 60 action classes, captured from three distinct camera views. It offers two evaluation protocols: Cross-Subject (X-Sub) and Cross-View (X-View). NTU-120 extends this to 120 classes and 114,480 samples, featuring Cross-Subject (X-Sub) and Cross-Setup (X-Set) protocols. PKU-MMD II encompasses 51 action classes; following standard practice, we utilize the skeleton modality under the Cross-Subject (X-Sub) protocol.
Focus on 3D skeleton. 2D skeleton sequences are typically extracted from images via pre-trained pose estimation networks [HRNet] and inherently represent coordinates on the image plane. Consequently, they can easily achieve high performance by being back-projected onto the image space to leverage existing image-domain representations [PoseC3D]. In contrast, 3D skeleton data encapsulates the pure geometric structure of human motion. Therefore, we focus our evaluation on 3D sequences to rigorously validate the model’s intrinsic geometric representation capability without any reliance on such image-domain priors.
@@ 4.2 Experiment Details
Architecture. To ensure a fair comparison with previous MAE-based methods [SkeletonMAE, MAMP, SJEPA, GFP], we employ a standard ViT [ViT] as our backbone encoder $g$. The encoder $g$ consists of 8 layers with a hidden dimension of $D=256$ and 8 attention heads, maintaining an identical architectural complexity to existing baselines. Input skeleton sequences are fixed to a temporal length of $T=64$ frames with $J=25$ joints. For patch embedding, we adopt a skeletal-temporal patch size of $P_{J}=1$ and $P_{T}=8$.
Pre-training setup. Our SLiM is pre-trained for 150 epochs within a teacher–student distillation framework [Teacher]. The teacher’s momentum $\tau$ is updated from 0.994 to 1. Optimization is performed using AdamW [AdamW] with a base learning rate of $2.0\times 10^{-4}$, incorporating a 20-epoch linear warmup followed by a cosine decay [CosineAnneal] down to $1.0\times 10^{-6}$. We train the model with a total batch size of 768 across 4 NVIDIA RTX A6000 GPUs.